9. A government agency has specialists who analyze the frequencies of letters of the alphabet in an attempt to decipher intercepted messages. In standard English​ text, a particular letter is used at a rate of 7.2​%

mu=

sigma=

b. In an intercepted​ message, a page of 2200 characters is found to have the letter occurring 272 times. Is this​ unusual?

A.Yes comma because 272 is greater than the maximum usual value.

B.No comma because 272 is within the range of usual values.

C.Yes, because 272 is within the range of usual values.

D.Yes comma because 272 is below the minimum usual value.
Answer #1

(a)mean(mu)=n*p=2200*0.072=158.2

standard deviation(sigma)= sqrt(n*p*(1-p))=sqrt(2200*0.072*(1-0.072)) =12.124

(b) Here

mean +/- 2*standard deviation =158.2 +/- 2*12.124=133.952,182.448

Since 272 is greater than the upper limit so this is unusual.

Correct option is (A)

A.Yes comma because 272 is greater than the maximum usual value.